# Emergence of Fixational and Saccadic Movements in a Multi-Level Recurrent Attention Model for Vision

### v1

@@ 5 Discussion
One of the key strengths of MRAM lies in its ability to produce emergent fixation and saccade patterns resembling human vision. MRAM takes inspiration from the human neural pathway of visual processing. As shown in Figures 2, 3, and 4, MRAM learns to alternate between short-range fixations for detail refinement and longer-range saccades for exploring new areas. These results align with Yarbus’ observations of human visual behavior [35], and statistics data of human eye movement [15], where fixations and long saccades emerge naturally during visual exploration. This behavior suggests that a hierarchical attention model can naturally approximate the dynamics of eye movement with repetitive fixation and exploratory saccade. Such patterns not only improve interpretability but also hint at the potential for MRAM to generalize well to other visual tasks that require more complex attention strategies. From a broader perspective, the experiment results reinforce the idea that the designed hierarchical attention mechanisms are not only computationally effective but also biologically plausible. The key limitations of MRAM lies in the unbalanced peripheral vision, that leads to the performance drop in the 2 scale setting, which mimic human visual system closer. The ablation study of CNN in DRAM suggested human brain have additional architecture to balance between recognition and gaze policy, that can be applied to more biological plausible hard attention model.
@@ 6.0.1 Acknowledgements
This study was funded by JST SPRING GX project (Grant Number JPMJSP2108).

### v2

@@ 5 Discussion
One of the key strengths of MRAM lies in its ability to produce emergent fixation and saccade patterns resembling human vision. MRAM takes inspiration from the human neural pathway of visual processing. As shown in Figures 2, 3, and 4, MRAM learns to alternate between short-range fixations for detail refinement and longer-range saccades for exploring new areas. These results align with Yarbus’ observations of human visual behavior [Yarbus], and statistics data of human eye movement [Saccadepattern1], where fixations and long saccades emerge naturally during visual exploration. This behavior suggests that a hierarchical attention model can naturally approximate the dynamics of eye movement with repetitive fixation and exploratory saccade. Such patterns not only improve interpretability but also hint at the potential for MRAM to generalize well to other visual tasks that require more complex attention strategies. From a broader perspective, the experiment results reinforce the idea that the designed hierarchical attention mechanisms are not only computationally effective but also biologically plausible. The key limitations of MRAM lies in the unbalanced peripheral vision, that leads to the performance drop in the 2 scale setting, which mimic human visual system closer. The ablation study of CNN in DRAM suggested human brain have additional architecture to balance between recognition and gaze policy, that can be applied to more biological plausible hard attention model.
@@ 6.0.1 Acknowledgements
This study was funded by JST SPRING GX project (Grant Number JPMJSP2108). This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections. The Version of Record of this contribution is published in Neural Information Processing. ICONIP 2025. Lecture Notes in Computer Science, vol 16310, and is available online at https://doi.org/10.1007/978-981-95-4378-6_21